OpenStudy (anonymous):

2+2? (:

OpenStudy (lukecrayonz):

5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

4 XD

OpenStudy (anonymous):

:3 NOOOOOOOOOO

OpenStudy (anonymous):

YESSSSSSSS XD XD XD

OpenStudy (anonymous):

2+2=4

OpenStudy (anonymous):

its 22 c; YAY

OpenStudy (anonymous):

how is it 22? ok I not even gonna say nothing

OpenStudy (anonymous):

I dont understand :F

OpenStudy (anonymous):

briszabo looks like a panda bear >:'3

OpenStudy (lukecrayonz):

x=y x^2=xy x^2-y^2=xy-y^2 (x+y)(x-y)=y(x-y) x+y=y 2y=y 2=1 1=0 Since 1=0 2+2+0=5

OpenStudy (skullpatrol):

Take 2 fingers in your left hand and put them with 2 fingers in your right hand, how many fingers are you showing?

OpenStudy (anonymous):

3 ^

OpenStudy (anonymous):

(:<

OpenStudy (anonymous):

Ok now you are clearly acting like your stupid

OpenStudy (anonymous):

HEY THANKS! ITS 3

OpenStudy (skullpatrol):

Can you count your fingers?

OpenStudy (anonymous):

Can you count your hair?

OpenStudy (anonymous):

(:<

OpenStudy (skullpatrol):

Yes, but it would take a very long time...

OpenStudy (anonymous):

WELL THEN EFF OFF SMARTY PANTS YAAAAYYYYYYYY

OpenStudy (skullpatrol):

Can you count clouds?

OpenStudy (anonymous):

maybe i can. MAYBE I CANT. ;)

OpenStudy (skullpatrol):

Yes, if they are too close to each other it is impossible, right?

OpenStudy (anonymous):

*GASP* SMARTY PANTS STRIKES AGAIN. c:

OpenStudy (skullpatrol):

Think about it, you don't have to be smart to understand...

OpenStudy (anonymous):

-CLAPS- BOOBS ARE GREAT.

OpenStudy (skullpatrol):

Can you count infinity?

OpenStudy (anonymous):

NOPE. CAN U ? ;)

OpenStudy (skullpatrol):

I can start but I'll never finish, right?

OpenStudy (anonymous):

GASP MAGIC

OpenStudy (anonymous):

this is funny

OpenStudy (anonymous):

LMAOOOOOO

OpenStudy (skullpatrol):

So, just like clouds, it DEPENDS on what and when you are counting 2 + 2

OpenStudy (anonymous):

DUUUUUUUUHHHHHHHHHHHHHHHOINK.

OpenStudy (anonymous):

o ^ o

OpenStudy (anonymous):

22! :D YAAAAYYYYYYYYY

OpenStudy (anonymous):

2 WOOTS FOR THE HOMO'S 2 WOOTS FOR THE BI'S 22 !!!! YAAAYYYYYYYYYY

OpenStudy (anonymous):

I DONT GIVE 2 BUTTS. :D PARTY IN THE USAAAA